Could you please help me?
I have a poster to create using the equation x^2 + 7x = -12
It is called Completing the Square
I have to show:
1. Set up the original equation to begin completing the square.
2. Show what is added to each side to complete the square (boxes)
3. Factor and evaluate the square roots
4. simplify any radicals and indicate plus/minus(_+)
5.Solve the resulting equations and indicate solutions (put in solution set braces)
Thank you so much

Answer by edjones(8007)

x^2 + 7x = -12
x^2+7x+ 49/4 = -12 + 49/4 completing the square.
(x + 7/2)^2=1/4
x= -7/2 + 1/2, x= -7/2 - 1/2
x=-3, x=-4
.
Ed